# Configuration spaces of circles in the plane

### v1

@@ 2.1. Configurations of disks in the plane
In this subsection we prove that configurations of disks in the plane encode the same information as configurations of points in the plane. We emphasize that we consider disks of all positive radii, anywhere in the plane. As a remark, if one only considers disks of a fixed radius, and restricts them to live in an infinite strip, then more interesting topology can occur [AKM21].
@@
$\Psi\colon\operatorname{LConf}_{n}(D^{2},\mathbb{R}^{2})\to\operatorname{LConf}_{n}(*,\mathbb{R}^{2})$
@@
$\overline{\Psi}\colon\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})\to\operatorname{Conf}_{n}(*,\mathbb{R}^{2})\text{.}$
Lemma 2.6.
@@
We first construct a map $\Phi$ in the other direction. Given $(\vec{x}_{1},\dots,\vec{x}_{n})$ in $\operatorname{LConf}_{n}(*,\mathbb{R}^{2})$, let $d>0$ be the minimum distance in $\mathbb{R}^{2}$ between the points in the tuple. For each $\vec{x}\in\mathbb{R}^{2}$, let $B_{d/3}(\vec{x})$ be the (closed) ball of radius $d/3$ centered at $\vec{x}$, so this is a disk. Now consider $(B_{d/3}(\vec{x}_{1}),\dots,B_{d/3}(\vec{x}_{n}))$. By the choice of $d$, these disks are pairwise disjoint, so this is an element of $\operatorname{LConf}_{n}(D^{2},\mathbb{R}^{2})$. Let
$\Phi\colon\operatorname{LConf}_{n}(*,\mathbb{R}^{2})\to\operatorname{LConf}_{n}(D^{2},\mathbb{R}^{2})$
@@
By construction, $\Psi\circ\Phi$ is the identity on $\operatorname{LConf}_{n}(*,\mathbb{R}^{2})$. In the other direction, $\Phi\circ\Psi$ sends a configuration of disks to the configuration of disks obtained by scaling each disk to have (its original center and) radius $d/3$, where $d$ is the minimum distance between centers of distinct disks. Thus the identity is (properly) homotopic to $\Phi\circ\Psi$, via a homotopy that first scales each disk with radius larger than $d/3$ down to have radius $d/3$, and then scales each disk with radius smaller than $d/3$ up to have radius $d/3$.
@@
The spaces $\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$ and $\operatorname{LConf}_{n}(D^{2},\mathbb{R}^{2})$ are connected, and have fundamental groups $B_{n}$ and $PB_{n}$ respectively. ∎
See Figure 6 for an example of a loop in $\operatorname{Conf}_{4}(D^{2},\mathbb{R}^{2})$, represented by a “braid of disks”. A feature of this picture is that it emphasizes how the disks in a configuration need not have the same radii.
Figure 6. A loop in $\operatorname{Conf}_{4}(D^{2},\mathbb{R}^{2})$, represented by a “braid of disks”.
@@ 3. Between circle configurations and trees
In contrast to $\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$, the configuration space $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ is disconnected for $n\geq 2$, as is $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$. Intuitively, a configuration in which two circles are nested cannot be continuously deformed to one in which no circles are nested; recall Figure 1 for two points in $\operatorname{LConf}_{2}(S^{1},\mathbb{R}^{2})$ that are in different components. For more circles, Figure 7 shows how a loop in $\operatorname{Conf}_{4}(S^{1},\mathbb{R}^{2})$ can be viewed as a braid, but where braids are partitioned off by larger enclosing circles. In this example, the largest circle represents the plane $\mathbb{R}^{2}$ and is not part of the configuration, two of the circles in the configuration are separately nested in another one, and the last circle in the configuration is un-nested with the others. The two circles nested in the other one may switch places, but no other permutations of circles are allowed. Circles of different radii are free to switch places, as long as the nesting pattern permits this. (To be clear, in the labeled case, no circles may switch places.)
Figure 7. A loop in $\operatorname{Conf}_{4}(S^{1},\mathbb{R}^{2})$, represented in a braid-like form.
@@
By a tree we will always mean a finite rooted ordered tree. In more detail: it is a finite directed graph that is connected and has no cycles (even undirected cycles). One vertex $\varnothing$ is designated as the root. The root has only outgoing edges, and every other vertex has one incoming edge and some number of outgoing edges. The vertices with zero outgoing edges are called leaves. Given a vertex $v$, the children of $v$ are the termini of the edges with origin $v$. The descendants of a vertex are its children, plus their children, etc. For each vertex $v$, we fix an ordering of the children of $v$, and view this as a “left-to-right” order for the purposes of drawing figures. A labeled tree is a tree with $n$ non-root vertices together with an identification of the non-root vertices with the set $\{1,\dots,n\}$. If we forget the labels on the vertices of a labeled tree, then we recover its underlying tree. We emphasize that the labels on the vertices need not have any relationship with the order on the children of a vertex.
There is some ambiguity about what it means for two trees or labeled trees to be equal versus isomorphic. Let us pin down exactly what we mean. When we refer to an isomorphism of trees, we mean this in the sense of abstract directed graphs, so an isomorphism between two trees is a bijection between their vertex sets that preserves adjacency and direction of edges, but need not preserve the order of any vertices. Note that there is only one self-isomorphism of a given tree that does preserve the order of each set of children of each vertex, namely the identity. Thus, any order-preserving isomorphism between two trees is necessarily unique, and so may be regarded as a notion of equality. For example, the trees in Figure 8 are isomorphic but not equal. (In this and all future figures, we will not draw arrows for directed edges in trees, but they are all directed downward. We will also draw the root larger than the other vertices.)
@@ 3.1. The labeled case
In this subsection we set up a connection between labeled configurations of circles and labeled trees. See [SW88, CCF+20] for other viewpoints of this same idea. Say that one circle in the plane is nested in another if the former lies in the bounded component of the complement of the latter. To be clear, we do not consider a circle to be nested in itself. Given a (labeled) configuration $\kappa=(C_{1},\dots,C_{n})$ of circles in the plane, if $C_{j}$ is nested in $C_{i}$ and moreover no $C_{k}$ simultaneously is nested in $C_{i}$ and has $C_{j}$ nested in it, then we will write $C_{i}\vdash C_{j}$. Also write $\mathbb{R}^{2}\vdash C_{i}$ whenever $C_{i}$ is not nested in any $C_{j}$. For $T$ a labeled tree with the non-root vertices labeled $1,\dots,n$, write $i\prec j$ whenever $j$ is a child of $i$, and write $\varnothing\prec i$ whenever $i$ is a child of the root $\varnothing$. As a remark, taking the transitive closure of $\prec$ yields a poset structure on $\{\varnothing,1,\dots,n\}$, with $T$ as its Hasse diagram (or rather $T$ drawn upside down, so the root is at the bottom, since in Hasse diagrams the partial order is directed upward).
Definition 3.2 (The labeled tree of a configuration).
@@
This gives us a canonical way of turning a labeled configuration of circles in the plane into a labeled tree. See Figure 10 for an example. In the other direction, we can turn labeled trees into configurations, although things are less canonical.
1111222233334444555566667777$\longrightarrow$1111222233334444555566667777
Figure 10. A labeled configuration $\kappa$ of circles in $\operatorname{LConf}_{7}(S^{1},\mathbb{R}^{2})$, and the associated labeled tree $T_{\kappa}$. Note that the left to right order on the children of a given vertex comes from the lexicographic order on the centers of the relevant circles.
Definition 3.3 (The fixed configuration for a labeled tree).
@@
$\varkappa_{T}=(\mathfrak{C}_{1}^{T},\dots,\mathfrak{C}_{n}^{T})$
of circles in the plane that satisfies the rule that $\mathfrak{C}_{i}^{T}\vdash\mathfrak{C}_{j}^{T}$ if and only if $i\prec j$. We construct $\varkappa_{T}$ as follows. First, write $i_{1},\dots,i_{k}$ for the labels on the children of the root, from left to right, and define the circle $\mathfrak{C}_{i_{j}}^{T}$ to have center $(\frac{j-1}{k},0)$ and radius $\frac{1}{3k}$. Note that these are pairwise disjoint, none are nested in each other, and they are all contained in the open disk of radius $1$. Now for any non-root vertex that is not a leaf, say with label $p$, write $i_{1},\dots,i_{k}$ for its children from left to right. Assume we have already constructed $\mathfrak{C}_{p}^{T}$, and that it is centered at $(x_{p},0)$ with radius $r_{p}$. Now for each $1\leq j\leq k$, construct $\mathfrak{C}_{i_{j}}^{T}$ to have center $(x_{p}+\frac{(j-1)r_{p}}{k},0)$ and radius $\frac{r_{p}}{3k}$. Note that these are pairwise disjoint, none are nested in each other, and they are all contained in $\mathfrak{C}_{p}^{T}$. Iterating this, we construct the entire tuple $\varkappa_{T}=(\mathfrak{C}_{1}^{T},\dots,\mathfrak{C}_{n}^{T})$, and get that it really is a (labeled) configuration of circles.
The root should be thought of as representing an unwritten 00th entry that is the unit circle. See Figure 11 for an example.
1111333344442222$\longrightarrow$1111333344442222
Figure 11. The labeled circle configuration $\varkappa_{T}$ corresponding to the given labeled tree $T$, constructed as in Definition 3.3. The circles labeled $4$ and $2$ are centered at $(0,0)$ and $(\frac{1}{2},0)$, and have radius $1/6$. The circles labeled $1$ and $3$ are centered at $(0,0)$ and $(\frac{1}{12},0)$, and have radius $1/36$.
Note that if the labeled trees $T$ and $U$ have the same underlying trees, then the underlying sets $\{\mathfrak{C}_{1}^{T},\dots,\mathfrak{C}_{n}^{T}\}$ and $\{\mathfrak{C}_{1}^{U},\dots,\mathfrak{C}_{n}^{U}\}$ are equal. Also note that the labeled tree of $\varkappa_{T}$ is $T$.
To be clear, we tend to use symbols like $\kappa$ and $C$ for arbitrary configurations and circles, and here we use $\varkappa$ and $\mathfrak{C}$ for these specific, fixed choices of configurations and circles.
@@
We should remark on the seemingly pathological case $n=0$. If $T$ is the trivial tree, that is, the tree with one vertex, then $\varkappa_{T}$ is the empty tuple, which is the only element of $\operatorname{LConf}_{0}(S^{1},\mathbb{R}^{2})=\{\emptyset\}$. In the other direction, the tree of the empty configuration consists only of a root, simply by definition. Thus, the $n=0$ case technically fits fine into all this setup, albeit in a somewhat non-intutive way.
@@ 3.2. The unlabeled case
The previous subsection was all about labeled configurations and labeled trees. In this subsection we establish the analogous connections between unlabeled configurations and unlabeled trees. The labeled case was easier to deal with first, bookkeeping-wise, and the unlabeled case follows more or less directly from it.
Recall from Definition 2.2 that $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ is the orbit space of $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ under the action of $S_{n}$. Write
$p\colon\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})\to\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$
@@
Let $\kappa=\{C_{1},\dots,C_{n}\}$ be an (unlabeled) configuration of $n$ circles. Let $\widetilde{\kappa}$ be any choice of labeled configuration whose underlying set is $\kappa$, e.g., we could use $\widetilde{\kappa}=(C_{1},\dots,C_{n})$, and consider the corresponding labeled tree $T_{\widetilde{\kappa}}$. Now the tree of $\kappa$, denoted $T_{\kappa}$, is defined to be the underlying tree of $T_{\widetilde{\kappa}}$. Since different labelings of $\kappa$ produce labeled trees with the same underlying tree, this is well defined.
Definition 3.6 (The fixed configuration for a tree).
@@
$\varkappa_{\widetilde{T}}=(\mathfrak{C}_{1}^{\widetilde{T}},\dots,\mathfrak{C}_{n}^{\widetilde{T}})$
from Definition 3.3. Let $\mathfrak{C}_{i}^{T}\coloneqq\mathfrak{C}_{i}^{\widetilde{T}}$ for each $i$, and set
$\varkappa_{T}\coloneqq\{\mathfrak{C}_{1}^{T},\dots,\mathfrak{C}_{n}^{T}\}\text{.}$
Note that the order of the $\mathfrak{C}_{i}^{T}$ depends on the choice of labeling, but as an (unordered) set, $\varkappa_{T}$ is well defined and equals the underlying set of $\varkappa_{\widetilde{T}}$. Also note that $p(\varkappa_{\widetilde{T}})=\varkappa_{T}$, and in fact the fiber of $\varkappa_{T}$ under $p$ equals the set of all $n!$ many points of the form $\varkappa_{\widetilde{T}}$ for each choice of labeled tree $\widetilde{T}$ with underlying tree $T$.

### v2

@@ 2.1. Configurations of disks in the plane
In this subsection we prove that configurations of disks in the plane encode the same information as configurations of points in the plane. This result is not new, see [Idr22, Lemma 5.49], but we include it for completeness. We emphasize that we consider disks of all positive radii, anywhere in the plane. As a remark, if one only considers disks of a fixed radius, and restricts them to live in an infinite strip, then more interesting topology can occur [AKM21].
@@
$\Psi\colon\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})\to\operatorname{Conf}_{n}(*,\mathbb{R}^{2})$
@@
$\overline{\Psi}\colon\operatorname{UConf}_{n}(D^{2},\mathbb{R}^{2})\to\operatorname{UConf}_{n}(*,\mathbb{R}^{2})\text{.}$
Lemma 2.6 (Lemma 5.49 of [Idr22]).
@@
We first construct a map $\Phi$ in the other direction. Given $(\vec{x}_{1},\dots,\vec{x}_{n})$ in $\operatorname{Conf}_{n}(*,\mathbb{R}^{2})$, let $d>0$ be the minimum distance in $\mathbb{R}^{2}$ between the points in the tuple. For each $\vec{x}\in\mathbb{R}^{2}$, let $B_{d/3}(\vec{x})$ be the (closed) ball of radius $d/3$ centered at $\vec{x}$, so this is a disk. Now consider $(B_{d/3}(\vec{x}_{1}),\dots,B_{d/3}(\vec{x}_{n}))$. By the choice of $d$, these disks are pairwise disjoint, so this is an element of $\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$. Let
$\Phi\colon\operatorname{Conf}_{n}(*,\mathbb{R}^{2})\to\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$
@@
By construction, $\Psi\circ\Phi$ is the identity on $\operatorname{Conf}_{n}(*,\mathbb{R}^{2})$. In the other direction, $\Phi\circ\Psi$ sends a configuration of disks to the configuration of disks obtained by scaling each disk to have (its original center and) radius $d/3$, where $d$ is the minimum distance between centers of distinct disks. Thus the identity is (properly) homotopic to $\Phi\circ\Psi$, via a homotopy that first scales each disk with radius larger than $d/3$ down to have radius $d/3$, and then scales each disk with radius smaller than $d/3$ up to have radius $d/3$.
@@
The spaces $\operatorname{UConf}_{n}(D^{2},\mathbb{R}^{2})$ and $\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$ are connected, and have fundamental groups $B_{n}$ and $PB_{n}$ respectively. ∎
See Figure 6 for an example of a loop in $\operatorname{UConf}_{4}(D^{2},\mathbb{R}^{2})$, represented by a “braid of disks”. A feature of this picture is that it emphasizes how the disks in a configuration need not have the same radii.
Figure 6. A loop in $\operatorname{UConf}_{4}(D^{2},\mathbb{R}^{2})$, represented by a “braid of disks”.
@@ 3. Between circle configurations and trees
In contrast to $\operatorname{UConf}_{n}(D^{2},\mathbb{R}^{2})$, the configuration space $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ is disconnected for $n\geq 2$, as is $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$. Intuitively, a configuration in which two circles are nested cannot be continuously deformed to one in which no circles are nested; recall Figure 1 for two points in $\operatorname{Conf}_{2}(S^{1},\mathbb{R}^{2})$ that are in different components. For more circles, Figure 7 shows how a loop in $\operatorname{UConf}_{4}(S^{1},\mathbb{R}^{2})$ can be viewed as a braid, but where braids are partitioned off by larger enclosing circles. In this example, the largest circle represents the plane $\mathbb{R}^{2}$ and is not part of the configuration, two of the circles in the configuration are separately nested in another one, and the last circle in the configuration is un-nested with the others. The two circles nested in the other one may switch places, but no other permutations of circles are allowed. Circles of different radii are free to switch places, as long as the nesting pattern permits this. (To be clear, in the labeled case, no circles may switch places.)
Figure 7. A loop in $\operatorname{UConf}_{4}(S^{1},\mathbb{R}^{2})$, represented in a braid-like form.
@@
By a tree we will always mean a finite rooted planar tree. In more detail: it is a finite directed graph that is connected and has no cycles (even undirected cycles). One vertex $\varnothing$ is designated as the root. The root has only outgoing edges, and every other vertex has one incoming edge and some number of outgoing edges. The vertices with zero outgoing edges are called leaves. Given a vertex $v$, the children of $v$ are the termini of the edges with origin $v$. The descendants of a vertex are its children, plus their children, etc. The tree is planar in the sense that we view it in the plane, with the root at the bottom and the edges directed upward. Planarity ensures that for each vertex $v$ there is an ordering of the children of $v$ from left to right. A labeled tree is a tree with $n$ non-root vertices together with an identification of the non-root vertices with the set $\{1,\dots,n\}$. If we forget the labels on the vertices of a labeled tree, then we recover its underlying tree. We emphasize that the labels on the vertices need not have any relationship with the planar order on the children of a vertex.
There is some potential ambiguity about what it means for two trees or labeled trees to be equal versus isomorphic. Let us pin down exactly what we mean. When we refer to an isomorphism of trees, we mean this in the sense of abstract directed graphs, so an isomorphism between two trees is a bijection between their vertex sets that preserves adjacency and direction of edges, but need not preserve the planar order of any vertices. Note that there is only one self-isomorphism of a given tree that does preserve the planar order of each set of children of each vertex, namely the identity. Thus, any isomorphism between two trees that preserves the planar order is necessarily unique, and so may be regarded as a notion of equality; in other words two (planar) trees are equal if they are isotopic in the plane. For example, the trees in Figure 8 are isomorphic but not equal. (In this and all future figures, we will not draw arrows for directed edges in trees, but they are all directed upward. We will also draw the root larger than the other vertices.)
@@ 3.1. The labeled case
In this subsection we set up a connection between labeled configurations of circles and labeled trees. See [SW88, CCF+20] for other viewpoints of this same idea. Say that one circle in the plane is nested in another if the former lies in the bounded component of the complement of the latter. To be clear, we do not consider a circle to be nested in itself. Given a (labeled) configuration $\kappa=(C_{1},\dots,C_{n})$ of circles in the plane, if $C_{j}$ is nested in $C_{i}$ and moreover no $C_{k}$ simultaneously is nested in $C_{i}$ and has $C_{j}$ nested in it, then we will write $C_{i}\vdash C_{j}$. Also write $\mathbb{R}^{2}\vdash C_{i}$ whenever $C_{i}$ is not nested in any $C_{j}$. For $T$ a labeled tree with the non-root vertices labeled $1,\dots,n$, write $i\prec j$ whenever $j$ is a child of $i$, and write $\varnothing\prec i$ whenever $i$ is a child of the root $\varnothing$. As a remark, taking the transitive closure of $\prec$ yields a poset structure on $\{\varnothing,1,\dots,n\}$, with $T$ as its Hasse diagram.
Construction 3.2 (The labeled tree of a configuration).
@@
This gives us a canonical way of turning a labeled configuration of circles in the plane into a labeled tree. See Figure 10 for an example.
$\longrightarrow11223344556677$
Figure 10. A labeled configuration $\kappa$ of circles in $\operatorname{Conf}_{7}(S^{1},\mathbb{R}^{2})$, and the associated labeled tree $T_{\kappa}$. Note that the left to right order on the children of a given vertex comes from the lexicographic order on the centers of the relevant circles.
In the other direction, we can turn labeled trees into configurations, although things are less canonical.
Construction 3.3 (The fixed configuration for a labeled tree).
@@
$\underline{\kappa}_{T}=(\underline{C}_{1}^{T},\dots,\underline{C}_{n}^{T})$
of circles in the plane that satisfies the rule that $\underline{C}_{i}^{T}\vdash\underline{C}_{j}^{T}$ if and only if $i\prec j$. We construct $\underline{\kappa}_{T}$ as follows. First, write $i_{1},\dots,i_{k}$ for the labels on the children of the root, from left to right, and define the circle $\underline{C}_{i_{j}}^{T}$ to have center $(\frac{j-1}{k},0)$ and radius $\frac{1}{3k}$. Note that these are pairwise disjoint, none are nested in each other, and they are all contained in the open disk of radius $1$. Now for any non-root vertex that is not a leaf, say with label $p$, write $i_{1},\dots,i_{k}$ for its children from left to right. Assume we have already constructed $\underline{C}_{p}^{T}$, and that it is centered at $(x_{p},0)$ with radius $r_{p}$. Now for each $1\leq j\leq k$, construct $\underline{C}_{i_{j}}^{T}$ to have center $(x_{p}+\frac{(j-1)r_{p}}{k},0)$ and radius $\frac{r_{p}}{3k}$. Note that these are pairwise disjoint, none are nested in each other, and they are all contained in $\underline{C}_{p}^{T}$. Iterating this, we construct the entire tuple $\underline{\kappa}_{T}=(\underline{C}_{1}^{T},\dots,\underline{C}_{n}^{T})$, and get that it really is a (labeled) configuration of circles.
The root should be thought of as representing an unwritten 0th entry that is the unit circle. See Figure 11 for an example.
$\longrightarrow11334422$
Figure 11. The labeled circle configuration $\underline{\kappa}_{T}$ corresponding to the given labeled tree $T$, as in Construction 3.3. The circles labeled $4$ and $2$ are centered at $(0,0)$ and $(\frac{1}{2},0)$, and have radius $1/6$. The circles labeled $1$ and $3$ are centered at $(0,0)$ and $(\frac{1}{12},0)$, and have radius $1/36$.
Note that if the labeled trees $T$ and $U$ have the same underlying trees, then the underlying sets $\{\underline{C}_{1}^{T},\dots,\underline{C}_{n}^{T}\}$ and $\{\underline{C}_{1}^{U},\dots,\underline{C}_{n}^{U}\}$ are equal. Also note that the labeled tree of $\underline{\kappa}_{T}$ is $T$.
To be clear, we tend to use symbols like $\kappa$ and $C$ for arbitrary configurations and circles, and here we use $\underline{\kappa}$ and $\underline{C}$ for these specific, fixed choices of configurations and circles.
@@
We should remark on the seemingly pathological case $n=0$. If $T$ is the trivial tree, that is, the tree with one vertex, then $\underline{\kappa}_{T}$ is the empty tuple, which is the only element of $\operatorname{Conf}_{0}(S^{1},\mathbb{R}^{2})=\{\emptyset\}$. In the other direction, the tree of the empty configuration consists only of a root, simply by definition. Thus, the $n=0$ case does fit into this setup, albeit in a somewhat non-intuitive way.
@@ 3.2. The unlabeled case
The previous subsection was all about labeled configurations and labeled trees. In this subsection we establish the analogous connections between unlabeled configurations and unlabeled trees. The labeled case was easier to deal with first, and the unlabeled case follows more or less directly from it.
Recall from Definition 2.2 that $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ is the orbit space of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ under the action of $S_{n}$. Write
$p\colon\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})\to\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$
@@
Let $\kappa=\{C_{1},\dots,C_{n}\}$ be an unlabeled configuration of $n$ circles. Let $\widetilde{\kappa}$ be any choice of labeled configuration whose underlying set is $\kappa$, e.g., we could use $\widetilde{\kappa}=(C_{1},\dots,C_{n})$, and consider the corresponding labeled tree $T_{\widetilde{\kappa}}$. Now the tree of $\kappa$, denoted $T_{\kappa}$, is defined to be the underlying tree of $T_{\widetilde{\kappa}}$. Since different labelings of $\kappa$ produce labeled trees with the same underlying tree, this is well defined.
Construction 3.6 (The fixed configuration for a tree).
@@
[…]
from Construction 3.3. Let $\underline{C}_{i}^{T}\coloneqq\underline{C}_{i}^{\widetilde{T}}$ for each $i$, and set
$\underline{\kappa}_{T}\coloneqq\{\underline{C}_{1}^{T},\dots,\underline{C}_{n}^{T}\}\text{.}$
Note that the order of the $\underline{C}_{i}^{T}$ depends on the choice of labeling, but as an (unordered) set, $\underline{\kappa}_{T}$ is well defined and equals the underlying set of $\underline{\kappa}_{\widetilde{T}}$. Also note that $p(\underline{\kappa}_{\widetilde{T}})=\underline{\kappa}_{T}$, and in fact the fiber of $\underline{\kappa}_{T}$ under $p$ equals the set of all $n!$ many points of the form $\underline{\kappa}_{\widetilde{T}}$ for each choice of labeled tree $\widetilde{T}$ with underlying tree $T$.